Systems Staff Meeting Notes 08/18/03

Systems Staff Meeting Notes
August 18, 2003

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. Cynthia’s dog had the pins removed from his leg. She forgot to bring them in to show everyone.
  2. Bobbie – the roof has been put on the new addition to her house.
  3. Will - Sea World was wetter than normal.
  4. Discussion: partly cloudy vs. partly sunny.
  5. Weekly moves meeting – construction status – Rich is currently staying with us in Systems. Due to the constant change in plans, the move meetings are now weekly and Bill has been asked to join. At the meeting last Friday, the VOIP cancellation was discussed: CNS phone did not want it, CNS nets did – they wanted the whole building. VOIP is back on again. Pricing model is in question. Data: $8 + phone: $16 = $24.25. The net crew wants to charge $8 on unused data line. They are currently rethinking how the dollars will flow. Nets would like the ability to assign VLAN IP on the phone. Geoffrey notes that people may try to use the ports anyways, if they are available. We can turn them off. The details are being worked out. This is scheduled to start October 1st. They will try to move people out of West in stages – phones, also. CIRCA access to be DHCP, details to be worked out. The phone system in East and MSL will be maxed out with the move. VOIP advantages include conference calls, interactive video. May switch to a new system and integrate everyone. East and Systems’ conference rooms will be made available to all. Systems loading dock will be parking for the construction crews. 22 employees will be working out at the DOT (Facilities, Access Services, mail, supplies). They have purchased another van; one will be stationed in the East loading dock. Window renovation in East is ongoing. Directors to occupy the 2nd, 3rd, and 4th floors of East – south tower. Personnel to occupy the 3rd floor tower. Room 1 to have 60 public computers, 16 people and printers. Circulation and reference desks to go in lobby (which currently has no power or phones). Suzy still relocated to DLC. 4th floor fiber pull – found key for rack, will move it to Systems and put switches and power supply in it.
Ongoing Systems Updates
  1. Some liaisons are trying to opt out of the program. They say there is not enough time to do the required work and their regular assignments.
  2. There is a bottleneck of old equipment not being returned. Much of it is needed in other places.
  1. The network is seemingly under attack this morning – special scanning of Unix, Samba & SSH is causing traffic.
  2. Service Pack 4 – still rolling out, slow going. Some people are having difficulty with the instructions. Reports of home computers getting whacked. The origin of the campus hit came form the dial-up pool.
  3. Trend scan – all machines should be at 618, but many are frozen at 600. Scans may have knocked the port out of synch. 54 subnet is up-to-date. Probable core patch is preventing those ports from updating. Need to contact Net Services about why it is not propagating.
Software Projects Status
  1. TILT request – online tutorial website for HSS& S. They are trying to adopt a system currently in place at the University of Texas. The request has not yet been approved.
  2. FilmLog changes finished.
  3. FCLA tests – Patron Loads – we’ll see if it can incorporate UF IDs this time.
  4. Aerial – DLC – Ying – not enough data yet.
  5. Cynthia is working on mods for SysRequest.
Web Projects
  1. Committee to meet later this week, then they’ll make a recommendation on Web master position.
DP Coordinator
  1. Fall semester starts next Monday. Students will be back on the desk.
  2. Nikia will work this week.

Back to the Systems Homepage

Last updated August 25, 2003
by Debra Fetzer